OpenStudy (dugalde):

Can you help me find the rectangular coordinates of each point (5,pi/4) (-2,pi/6) (-1,2pi/3)

jimthompson5910 (jim_thompson5910):

I'm assuming each point is in polar form (r,theta)

OpenStudy (dugalde):

Yes they are

jimthompson5910 (jim_thompson5910):

You can use these formulas x = r*cos(theta) y = r*sin(theta) make sure you are in radian mode

OpenStudy (dugalde):

so 5*cos(theta) and pi/4*sin(theta)

jimthompson5910 (jim_thompson5910):

(r,theta) = (5,pi/4) so r = 5 theta = pi/4

jimthompson5910 (jim_thompson5910):

x = r*cos(theta) = 5*cos(pi/4) = ?? y = r*sin(theta) = 5*sin(pi/4) = ??

OpenStudy (dugalde):

(3.54, 3.54)

OpenStudy (jdoe0001):

\(\large \begin{array}{cccllll} 5&,&(\frac{\pi}{4})\\ r&&\theta \end{array}\qquad \begin{cases} x=rcos(\theta)\\ y=rsin(\theta) \end{cases}\qquad \begin{array}{cccllll} x&,& y\\ \uparrow &&\uparrow \\ rcos(\theta)&&rsin(\theta) \end{array}\)

jimthompson5910 (jim_thompson5910):

since \[\Large \cos\left(\frac{\pi}{4}\right) = \frac{\sqrt{2}}{2}\] and \[\Large \sin\left(\frac{\pi}{4}\right) = \frac{\sqrt{2}}{2}\] this means \[\Large (x,y) = \left(\frac{5\sqrt{2}}{2},\frac{5\sqrt{2}}{2}\right)\]

jimthompson5910 (jim_thompson5910):

that is the exact form, but the approximate form works as well

OpenStudy (dugalde):

ok thanks!

jimthompson5910 (jim_thompson5910):

np